# What Would an LLM Do? Evaluating Policymaking Capabilities of Large Language Models

### v1

@@ Introduction
LLM as a Policymaker
Our homelessness mitigation benchmark is presented to LLMs, which are prompted to respond as decision makers representing a non-profit organisation. Here we describe our choice of LLMs and baseline prompts.
@@
We experiment with the following LLMs, listed in increasing order of the number of model parameters (which are shown in parentheses): Granite 3.1 8B Instruct (8 B), GPT 4.1 Mini (8 B), Mixtral 8X7B Instruct (around 47 B), Llama 3.3 70 B Instruct (70 B), Deepseek V3 (671 B), GPT 4.1 (estimated around 1.8 T). These models capture a broad spectrum of characteristics, spanning a range of sizes and architectures – from compact, efficiency-optimized models (e.g., Granite 3.1, GPT 4.1 Mini) to larger frontier models (e.g., Deepseek V3, GPT 4.1). This diversity enables comparisons across performance tiers, supporting robust conclusions about decision quality, consistency, and ethical alignment.
@@
We prompt LLMs in two primary ways, one for each of the two tasks. In one approach, we ask the LLM to select one policy among the $4$ choices for scenarios in the benchmark. In another approach, we ask the LLM to provide an ordered ranking of their choices among the policies. In both cases, we also request the LLM to provide a brief justification for their response. For the GPT models, we employ a constrained reasoning prompt that guides models through four steps—summarizing the dilemma, identifying capabilities restored by each policy, assessing pros and cons, and outputting a ranking with justification. To encourage transparency without eliciting verbose chain–of–thought, we limit reasoning to under 150 characters per step. In addition, we conduct some experiments with ablation studies where LLMs are asked to make choices using additional considerations, such as paying heed to contextual information. We refer the reader to Appendix C where we provide some example prompts and prompt templates.
Assessing the Social Effects of Homelessness Mitigation Policies
We explore the social impact of the selected policies in an existing agent-model simulation (Aguilera et al. 2024) contextualized in the city of Barcelona, which is designed to evaluate the impact of homelessness alleviation policies.
@@
In the pipeline, LLMs convert the natural language policy proposals they have previously selected into agents’ behavioral adjustments in the ABM. This is done by requesting the LLM to update the state-action-transition (SAT) matrix accordingly, which regulates the behavior of agents during the simulation. Due to space limitations, we relegate details about the mathematical foundations of the ABM pipeline to Appendix B. This approach enables us to test the effects of a policy in terms of needs satisfaction of the agents representing people experiencing homelessness (PEH).
@@
We prompt an LLM by giving a specific role of a public policy expert and providing the following information: a description of the SAT matrix structure, the policy to be tested, a step-by-step description on how the matrix can be technically modified and the indication that the policies are framed in the Capability Approach. The LLM produces an update of the SAT matrix, modifying the behavior of agents in the simulation. More details on the specific prompt can be found in Appendix C.
Empirical Investigation
@@
Figure 2: Pairwise comparison of the choices and judgments of various LLMs using heat maps of the following: a) the fraction of common top choices, b) similarity between justifications for top choices (as measured by the ROUGE-L metric), and c) comparison of rankings of choices (as measured by the normalized Kendall tau distance).
Investigating LLMs’ Choices
@@
We explore how the various LLMs compare in terms of their choices and judgments on the homelessness mitigation benchmark. Each LLM is asked to consider all $160$ contextualized decision scenarios – $40$ each across the $4$ locations – and make a judgment about their top choice and rankings. Figure 2 shows pairwise comparison plots between all $6$ LLMs under consideration in the form of heat maps, using different aggregate metrics across panels. Panel (a) displays the fraction of scenarios where the top choices for a pair of models are identical. Panel (b) shows the average similarity between the textual justifications of LLMs for their top choice, as measured by the ROUGE-L similarity metric. Panel (c) compares rankings of choices, as gauged by the normalized Kendall tau distance, which is $1$ for identical rankings and 1-1 for completely reversed rankings.
@@ Results.
The plots in Figure 2 indicate that LLMs could potentially disagree about their top choices, particularly those from different model families. Panel (a) highlights that the smaller Granite 3.1 8B has less overlapping top choices, while other models choose the same policy around $70-80\%$ of the time. We also note that the ROUGE-L scores in panel (b) are substantially lower than the range of 0.40.50.4-0.5 (often regarded to represent moderate similarity), indicating that neutral prompts result in different justifications across LLMs on the benchmark. While there are well-known limitations of using ROUGE-L as a metric for comparing text, it remains popular for closely related tasks such as summarization. A comparison of rankings in panel (c) reflects similar trends. Overall, the results indicate that the choice of LLM affects policy choices and judgments on the benchmark.
Comparing Choices of LLMs vs. Experts
@@
Our comparison with experts operates on two levels: agreement on the single top-ranked policy and similarity across the entire ranking of four policies, measured using the normalized Kendall tau distance. This allows us to assess not only the final choice but also the structure of preference. The comparison between the two human experts on Universal scenarios provides a baseline for inter-human agreement on decontextualized policy dilemmas.
@@
Table 1: Alignment b/w LLMs vs. experts: the low correlation for the South Bend pair, despite high top-choice agreement, suggests a divergence in underlying policy priorities.
@@
   GPT-4.1 vs. JHB Expert 50% 0.20
   GPT-4.1 vs. SB Expert 60% 0.10
@@
   Human Baseline 40% 0.07
@@
Our analysis reveals a moderate but nuanced alignment between LLMs and experts. As shown in Table 1, which compares experts with GPT-4.1, the degree of agreement varies by context. In Johannesburg, the LLM’s top choice matched the expert’s choice 50% of the time, with a weak ranking correlation (Kendall’s Tau) of 0.20. The South Bend case is particularly revealing: despite a higher top-choice agreement of 60%, the overall ranking correlation was a mere 0.10. This suggests that while the LLM and the South Bend expert could often agree on the single best policy, their underlying priorities for the remaining options diverged significantly.
The Universal scenarios provide the most telling baseline. The two human experts that annotate the universal scenarios agreed on a top policy in only 40% of cases, and their overall rankings showed almost no correlation ($\tau=0.07$). In this light, GPT-4.1’s alignment with the South Bend expert (60% agreement, $\tau=0.33$) is notable, as it exceeds the human-human baseline on both metrics. We conjecture that this is because the experts are taking contextual information into account while making their choice, which can be vital for effective policy making.
Figure 4: Comparing top choice overlap between 4 LLMs and an expert in Johannesburg, with and without prompting LLMs to pay attention to the context when choosing.
We investigate this further in Figure 4, where we consider scenarios in Johannesburg while comparing top choices of the four non-GPT LLMs with those from the expert in four ways: over the 10 contextualized (Local) scenarios and the 10 Universal scenarios, both with and without additional prompting to the LLM to explicitly consider the local geographical context. We observe that in general: 1) there is more expert choice overlap in Local rather than Universal settings, and 2) specifying the context impacts LLMs’ choices and yields more overlap, even when scenarios are Local, i.e., already somewhat contextualized by design.
Beyond pure agreement rates, we find systematic divergence in the types of capabilities prioritized, as visualized in Figure 3. Experts’ top choices frequently centered on policies restoring practical reason (e.g., skills training, participatory governance) and affiliation (e.g., peer support, community integration). In contrast, GPT-4.1 demonstrated a stronger tendency to select policies that protect against immediate physical threats, prioritizing life, bodily health, and bodily integrity. The LLM appears to exhibit a form of heightened risk aversion focused on fundamental safety, whereas experts more readily balance safety with long-term empowerment and social cohesion. For instance, in the scenario titled “Safe-Parking for Vehicle Dwellers” (South Bend), the human expert preferred ‘Conversion-to-Housing Incentive Grants‘ to transition residents into stable rooms, emphasizing long-term capability expansion. GPT-4.1 favored a ‘Monitored Safe-Parking Lot‘, echoing its bias toward scalable, immediate hazard reduction.
This value divergence is starkly illustrated in a Johannesburg scenario regarding an open-air drug scene. The expert prioritized a “Broken Windows” policing surge (Capability: bodily integrity via public order). GPT-4.1 ranked this last, instead choosing a supervised consumption site (Capabilities: life, bodily health). Its verbal justification noted that policing “risks rights for limited gain,” while the site “saves most lives”. This highlights a fundamental difference: the expert chose an enforcement-based solution to restore order, while the LLM opted for a health-based, harm-reduction solution to prevent death. In summary, while an LLM’s choices could potentially be as consistent with an expert as that expert is with another, its underlying decision-making calculus suggests a discernible bias towards scalable, safety-first interventions over more relational or structural ones.
Gauging the Impact of Policies using ABMs
@@
We evaluate the concrete impacts of some of the benchmarked policies in terms of whether they satisfy the needs of the agents (representing PEH) as described in Section Assessing the Social Effects of Homelessness Mitigation Policies. We run simulations with 80 agents for 1450 steps (corresponding to a time frame of 2 months in the model), and use Deepseekr1 (with a temperature of 0.1) to update the SAT matrix on the scenarios where the expert top policy prioritization choice differs from the GPT4.1 top policy prioritization for the city of Barcelona (these are scenarios 1, 3 and 5 in the benchmark). The ABM framework used in this experiment is specifically designed for the city of Barcelona at the moment and will be extended further in future work (please see Appendix B for details). We run 10 simulations without the behavioral changes from the applied policy (with the base SAT matrix of the model) and 10 simulations with the behavioral changes induced by the policy (with the LLM-updated SAT matrix). This approach provides us with insights on how the policies selected by the expert and the LLMs differ in terms of the satisfaction of the needs of the agents (representing PEH).
@@
Table 2 summarizes the changes observed in the ABM simulation when applying either the LLM or the expert policy. For each category of needs, ‘Mean’ represents the difference of the average satisfaction value for PEH agents with or without the policy. ‘Std’ represents the difference in the dispersion of needs satisfaction with or without policies. And finally, ‘p-value’ is associated with a standard t-test to assess whether or not the changes in terms of needs’ satisfaction were statistically significant.
@@
Based on the results of Table 2, the LLM-recommended policies seem to demonstrate superior overall balance compared to the expert-recommended policies. While both approaches notably enhanced safety and self-esteem needs, the improvements from LLM-implemented policies were more substantial. Additionally, LLM implementations consistently prevented the negative effects on belonging needs that were seen with expert policies in scenarios 1 and 5. Both policy types successfully reduced outcome variability (negative change in std values), but LLMs achieved this while maintaining positive or neutral effects across all need categories—a balance experts couldn’t sustain. More detailed results are available in Appendix D.
Discussion and Conclusions
This paper introduces a novel benchmark to evaluate whether LLMs are aligned with human domain experts to inform social policymaking, in particular on the topic of homelessness alleviation. The benchmark includes 170 homelessness-related policy scenarios, grounded in the Capability Approach for human development, across four geographic contexts. It further presents an automated pipeline that links LLM-generated policy proposals to an agent-based modelling framework. This enables to explore the social impact of the recommended policies in a non-invasive manner through simulated social scenarios. The article opens a path to leverage LLMs for social policy making, providing humans with alternative, contextualized policies at scale, and an indication of their social impact.
@@
The empirical investigation unveils that LLMs policy recommendation patterns seem to apply a highly stable internal heuristic, i.e. prioritize immediate physical safety and broad coverage—across different contexts. While this leads to dependable internal logic, it can also expose a context-blind rigidity. Experts, in contrast, demonstrably tailored their decisions to address locally encoded sociopolitical realities that can aggravate the ostracism faced by PEHs, such as related to ethnicity (in South Africa) or to religious minorities (in South Bend). Moreover, our analysis reveals a moderate higher alignment on LLMs’ top choice with the two experts located in the Global North. However, when asking LLMs to focus in a contextualized scenario, the results become more closely aligned with the local expert (as we describe for Johannesburg). Our findings also compare the social impact of the policies selected by LLMs vs those selected by the domain expert in the city of Barcelona, in an agent-based modelling framework. The results obtained in the simulated social scenario show that LLMs recommended policies could be better fulfilling the overall needs of PEH population.
As in most computational approaches to complex social topics, the paper contains some limitations that will be addressed in future work. First of all, policies were manually annotated by only one domain expert per location. Although these domain experts work in close collaboration with specialized local non-profits on homelessness alleviation, feedback from additional experts will provide a broader representation of the local context. Secondly, our study has only been conducted with LLMs in English, which is not the most spoken language in all the locations in scope. Multilingual analysis will be included in future analysis. Finally, the agent-based modelling framework only handles scenarios for one of the locations in scope (Barcelona) and is only able to simulate some of the policies in scope. Future work will expand the simulated scenarios to obtain more comprehensive information on the social impact of the proposed policies.
These results reveal promising potential for LLMs to provide new insights on social policy making, if the analysis is rooted in a conceptual ethical framework which is relevant to the social topic in scope – our study is framed in the Capability Approach for human development – and when responsible guardrails and contextual calibrations are introduced in collaboration with local domain experts.

### v2

@@ 3.1.5 Model Evaluation Tasks
3.2 LLMs as Policymakers
Our homelessness mitigation benchmark is presented to LLMs, which are prompted to respond as decision makers representing a non-profit organization. Here we describe our choice of LLMs and baseline prompts.
@@ 3.2 LLMs as Policymakers
We experiment with the following LLMs, listed in increasing order of number of model parameters: Granite 3.1 8B Instruct (8 B), GPT 4.1 Mini (8 B), Mixtral 8X7B Instruct ($\sim$ 47 B), Llama 3.3 70 B Instruct (70 B), Deepseek V3 (671 B), GPT 4.1 (estimated $\sim$ 1.8 T). These models capture a broad spectrum of characteristics, spanning a range of sizes and architectures, enabling richer comparisons.
@@
We prompt LLMs in two primary ways, one for each task. In one approach, we ask the LLM to select one policy among the $4$ choices per scenario; in the second, we ask the LLM to rank their choices. In both cases, we also request the LLM to provide a brief justification for their response. For the GPT models, we employ a constrained reasoning prompt that guides models through four steps: summarizing the dilemma, identifying capabilities restored by each policy, assessing pros and cons, and outputting a ranking with justification. In addition, we conduct some experiments with ablation studies where LLMs are asked to make choices using additional considerations, such as paying heed to contextual information. We refer the reader to Appendix C where we provide example prompts and prompt templates.
3.3 Assessing Social Impact with ABMs
We evaluate the social impact of policies chosen by LLMs vs. those proposed by human domain experts in the city of Barcelona, using an existing agent-model simulation which is purposely built to simulate homelessness policymaking in this city Aguilera et al. (2024).
@@ 3.3 Assessing Social Impact with ABMs
We present a pipeline that uses LLMs as structured policy recommenders. In this pipeline, LLMs convert the natural language policy proposals they have previously selected into agents’ behavioral adjustments in the ABM. This is done by requesting the LLM to update the state-action-transition (SAT) matrix accordingly, which regulates the behavior of agents during the simulation. Due to space limitations, we relegate details about the mathematical foundations of the ABM pipeline to Appendix B. This approach enables us to compare the effects of policies generated by LLMs and by domain experts in terms of needs satisfaction of the agents representing people experiencing homelessness (PEH).
@@
We prompt an LLM by giving a specific role of a public policy expert and providing the following information: a description of the SAT matrix structure, the policy to be evaluated, a step-by-step description on how the matrix can be technically modified and the indication that the policies are framed in the Capability Approach. The LLM produces an update of the SAT matrix, modifying the behavior of agents in the simulation. Prompt details are in Appendix C.
4 Empirical Investigation
@@
Figure 2: Pairwise comparison of the choices and judgments of various LLMs using heat maps of the following: a) the fraction of common top choices, b) similarity between justifications for top choices (as measured using Sentence-BERT embeddings), and c) comparison of rankings of choices (as measured by the normalized Kendall tau distance).
4.1 Investigating LLMs’ Choices
@@
We explore how the various LLMs compare in terms of their choices on the homelessness mitigation benchmark. Each LLM is asked to consider all $160$ contextualized decision scenarios ($40$ each across $4$ locations) and make a judgment about their top choice and rankings using neutral prompts. It is important to note that there is no circularity in this evaluation – while LLMs assisted in generating policy options for each scenario, the ranking task is fundamentally different: models must reveal their preferences among a fixed set of policies, thereby indirectly revealing how they prioritize different capabilities within the Capability Approach framework. Figure 2 shows pairwise comparison plots between all $6$ LLMs under consideration in the form of heat maps, using different aggregate metrics across panels. Panel (a) displays the fraction of scenarios where the top choices for a pair of models are identical. Panel (b) shows the average semantic similarity between the textual justifications of LLMs for their top choice, as measured using Sentence-BERT embeddings. Panel (c) compares rankings of choices, as gauged by the normalized Kendall tau distance, which is $1$ for identical rankings and 1-1 for completely reversed rankings.
@@ Results.
The plots in Figure 2 indicate that there may be disagreements between LLMs, particularly those from different model families. Panel (a) highlights that the smaller Granite 3.1 8B has less overlapping top choices, while other models choose the same policy around $70-80\%$ of the time. Panel(b) suggests that semantic similarity between justifications is notably high across all pairs (0.71-0.84), indicating convergence in reasoning patterns and linguistic styles. However, normalized Kendall tau correlations between full rankings are considerably lower (0.31-0.65) in panel (c), suggesting models use similar justification frameworks but apply different weightings when ordering options. These findings reveal that models agree most on how they justify, less on what they choose first, and least on how they rank the policies – perhaps because surface-level homogenization in articulation masks diverse underlying preferences.
4.2 Comparing Choices of LLMs vs. Experts
@@
Our comparison with experts operates on two levels: agreement on the top-ranked policy choice as well as the ranking over four policies, measured using the normalized Kendall tau distance. The comparison between the human experts on Universal scenarios provides a reference for inter-human agreement on decontextualized policy dilemmas.
@@
Table 1: Alignment Scores: GPT-4.1 vs. Experts. Location-specific results represent the average alignment across multiple domain experts. Inter-Expert Consensus rows indicate the agreement between the human experts themselves.
@@
     GPT-4.1 vs. BCN Experts (Avg.) 50% 0.32
        BCN Inter-Expert Consensus 50% 0.23
     GPT-4.1 vs. JHB Experts (Avg.) 55% 0.28
        JHB Inter-Expert Consensus 50% 0.50
     GPT-4.1 vs. SB Experts (Avg.) 50% 0.22
        SB Inter-Expert Consensus 20% 0.17
@@
     GPT-4.1 vs. Exp. 3 (BCN) 40%
     Human Inter-Expert Consensus 33% 0.07
@@
First, we compare GPT-4.1 with multiple experts across urban contexts. As shown in Table 1, alignment varies by city but frequently meets or exceeds inter-expert consensus. In Johannesburg, the LLM’s top choice matched expert priorities in 55% of cases ($\tau=0.28$), exceeding the 50% consensus between experts. In South Bend, the model’s 50% top-choice agreement ($\tau=0.22$) also outperformed the 20% consensus between experts. Barcelona showed 50% alignment ($\tau=0.32$), matching the consensus between experts.
The Universal scenarios provides a critical reference point for these findings. Comparing three experts across regions shows a low average consensus of 33% ($\tau=0.07$), reflecting the importance of local context in policy prioritization. In this light, GPT-4.1’s strong alignment with the South Bend expert on these universal dilemmas (60% agreement, $\tau=0.33$) is notable. This disparity could potentially be illustrating a bias, possibly stemming from the over-representation of US-centric published discourse in their training data. Conversely, the lower agreement with experts from Johannesburg (20%) and Barcelona (40%) on identical universal dilemmas may indicate a gap in the model’s ability to align with the local contexts, regarding the topic of homelessness policymaking, in geographic locations from where there is less training data in frontier models. However, this pattern should not be over-interpreted. Differences between experts’ opinions may arise due to heterogeneity in their underlying policy philosophies, or due to the way Universal scenarios are contextualized. We therefore treat this as a hypothesis that motivates further investigation rather than a claim about the nature of the training-data composition.
Figure 4: Comparing top choice overlap between 4 LLMs and the primary domain expert in Johannesburg, with and without prompting LLMs to consider the local context when selecting policies.
@@
We further investigate the importance of contextualization in Figure 4, where we consider the scenarios in Johannesburg, comparing top choices of the four non-GPT LLMs with those from the primary expert in four ways: over the 10 contextualized (Local) scenarios and the 10 Universal scenarios, both with and without additional prompting to the LLM to explicitly consider the local geographical context. We observe that: 1) there is higher alignment with the domain expert choice in the Local rather than Universal settings, and 2) specifying the context impacts LLMs’ choices and yields higher alignment, even when scenarios are Local, i.e., already somewhat contextualized by design.
Beyond policy choice agreement rates, we find systematic divergence in the types of capabilities prioritized, as visualized in Figure 3. Experts’ top choices frequently centered on policies restoring practical reason (e.g., skills training, participatory governance) and affiliation (e.g., peer support, community integration). In contrast, GPT-4.1 demonstrated a stronger tendency to select policies that protect against immediate physical threats, prioritizing life, bodily health, and bodily integrity. For instance, in the scenario titled “Safe-Parking for Vehicle Dwellers” (South Bend), the human expert preferred ‘Conversion-to-Housing Incentive Grants‘ to transition residents into stable rooms, emphasizing long-term capability expansion. GPT-4.1 favored a ‘Monitored Safe-Parking Lot‘, echoing its preference toward scalable, immediate hazard reduction. Further studies should be conducted to identify if LLMs’ preferences pertaining to the prioritization of safety-first interventions over relational or structural ones is connected to the over-representation of data from specific parts of the world and the public narrative on homelessness in these regions.
4.3 Gauging the Impact of Policies using ABMs
@@
We evaluate benchmarked policies by measuring needs satisfaction for agents representing persons experiencing homelessness (PEH) within our ABM pipeline. Simulations use 80 agents over 1,450 steps (simulating two months). We employ Deepseek-R1 ($T=0.1$) to update the SAT matrix for Barcelona-specific scenarios where expert and GPT-4.1 policy prioritizations diverge (Scenarios 1, 3, and 5). While this framework is currently tailored to Barcelona, future work will include three additional cities (see Appendix B). We compare 10 baseline simulations using the original SAT matrix against 10 simulations using the LLM-updated matrix. This approach highlights how expert-selected versus LLM-suggested policies differ in addressing PEH needs satisfaction.
@@ Results.
Table 2 summarizes the changes observed in the ABM simulation when applying either the policies recommended by LLMs or by the domain experts. For each category of needs, ‘Mean’ represents the difference of the average satisfaction value for PEH agents with or without the policy. ‘Std’ represents the difference in the dispersion of needs satisfaction with or without policies. And finally, ‘p-value’ is associated with a standard t-test to assess whether the changes in terms of needs’ satisfaction were statistically significant.
@@
Based on results in Table 2, LLM-recommended policies seem to demonstrate a slightly superior overall balance compared to expert-recommended ones. While both approaches notably enhance safety and self-esteem needs, improvements from LLM choices appear to be more substantial. Both policy types successfully reduced outcome variability (negative change in std values), but LLMs achieved this while maintaining positive or neutral effects across all need categories. More detailed results are available in Appendix D.
5 Discussion and Conclusion
This paper investigates the potential of LLMs to support policymaking on the subject of homelessness alleviation in four global cities, using a benchmark grounded in the Capability Approach for human development and supported by an agent-based simulation. The study is motivated by the expressed need of government officials, domain experts and non-profit organizations, who need to make informed decisions in a complex and evolving environment, often with limited time and data. The capacity of LLMs to leverage vast amounts of humanly-generated existing data and explore flexible role-play scenarios (i.e. acting as domain experts) make them uniquely suited to potentially provide new insights for the complexity of social policymaking. However, given the current limitations of LLMs, in terms of transparency, explainability, security and ethics, among others, there are open questions around whether policies recommended by LLMs are to be trusted and can be helpful for decision makers.
@@ 5 Discussion and Conclusion
Our empirical investigation about LLM policy recommendations suggests that while models may have learned similar ways to articulate reasoning, they maintain diversity in how they weight and order the options. Importantly, LLMs’ choices seem to prioritize immediate physical safety across the different contexts, over policies that aim to improve the social affiliation of the affected population and mitigate ostracism. This could be exposing a context-blind rigidity and a potential bias based on the published online discourse on homelessness in the regions of the world with over-represented data in LLMs. Experts, in contrast, demonstrably tailored their decisions to address locally encoded sociopolitical realities that can aggravate the ostracism faced by PEHs, such as related to ethnicity (in South Africa) or to religious minorities (in South Bend). However, for a selection of the scenarios in the benchmark simulated in the ABM, our findings show that the social impact of policies selected by LLMs slightly better fulfill the overall needs of PEH population, vs. those selected by the domain expert in Barcelona.
As is the case in most computational approaches to complex social topics, our work has several limitations that need to be addressed in future work. First of all, a wider diversity of non-profit organizations and domain experts should be included in future policy scenario creation and prioritization. Even though the baseline scenario was created in close collaboration with domain experts, the remaining scenarios in the benchmark rely heavily on LLM assistance, albeit with detailed designed prompting that incorporates the Capability Approach and the local contexts. Furthermore, benchmark scenarios and policies are currently all in English. We recommend that future studies build the entire benchmark (not only the baseline) with realistic policies proposed by domain experts in the local languages of the cities in scope. Finally, the agent-based modeling framework that allows one to evaluate the impact of the policies suggested by LLMs, versus those recommended by domain experts, currently handles the policy scenarios for one of the locations in scope only (Barcelona). The pipeline in between LLM generated policies and the ABM framework aims to open up a line of work to evaluate LLM-suggested policies through ABM frameworks in the different locations in scope at scale.
Our current findings highlight both promise and caution for the use of LLMs to inform social policymaking. On the one hand, LLMs show a certain degree of alignment with domain experts and propose plausible policy recommendations, which can be valuable inputs to policy makers and non-profit organizations in the context of limited data and time constraints. On the other, our findings illustrate that LLMs are less sensitive to local contexts than domain experts, which could be influenced by the over-representation of online data from certain parts of the world. In addition, LLMs seem to not sufficiently consider the restoration of central human development capabilities (such as the need for affiliation) as compared to domain experts, who emphasized measures dedicated to mitigate social ostracism against PEHs. Moving forward, the responsible use of of LLMs in social policymaking contexts will require not only technical refinement, but deliberate mechanisms of evaluation in line with conceptual ethical frameworks, such as the Capability Approach of human development, as well as responsiveness to local contexts and valuable inputs from domain experts.